Would anyone please help me write the slope please...
Write the slope-intercept form of the line through (0, -2) that is parallel to the line 5x + 4y =3

Found 2 solutions by stanbon, JimMarshall:
Answer by stanbon(75887)

The slope of the given line is m=-5/4
You want a line of the form y=mx+b
You have y=-2, x=0, m=-5/4
You need to solve for "b"
Substituting into the form we can solve for "b" as follows:
-2=(-5/4)(0)=b
b=-2
The equation you want is y=(-5/4)x-2
Cheers,
Stan H.

Answer by JimMarshall(82)

The slope of the given line is m=-5/4
You want a line of the form y=mx+b
You have y=-2, x=0, m=-5/4
You need to solve for "b"
Substituting into the form we can solve for "b" as follows:
-2=(-5/4)(0)=b
b=-2
The equation you want is y=(-5/4)x-2 Ans